what is the spread between short and long term yield for 4 zero coupons with $10000 face value with the ff term and price.

term price

1 yr 9700.87

2 yr 9333.51

3 yr 8880.99

4 yr 8220.70

how can the curve be described?
Answer #1

1 yr 9700.87

2 yr 9333.51

3 yr 8880.99

4 yr 8220.70

Rate of a zero coupon bond = [(Face value/current price)^(1/n)]-1
n= number of years to maturity.

Rate 1 year = [(10,000/9700.87)^(1/1)]-1= 3.084%

Rate 2 year = [(10,000/9333.51)^(1/2)]-1= 3.509%

Rate 3 year = [(10,000/8880.99)^(1/3)]-1= 4.035%

Rate 4 year = [(10,000/8220.70)^(1/4)]-1= 5.02%

the spread between short and long term yield is 5.02%-3.084% = 1.936%

The yield curve is upward sloping, which is very common.